OpenStudy (anonymous):

-4^2 +2^-1 Express the number in a/b where a and b are integers

OpenStudy (anonymous):

\[\left( -4 \right)^{2}+2^{-1}=16+\frac{ 1 }{ 2 }=\frac{ 32+1 }{ 2 }=\frac{ 33 }{2 }\]

OpenStudy (anonymous):

how did you get 1/2 and how did you get 32

OpenStudy (anonymous):

that answer is wrong

OpenStudy (anonymous):

\[2^{-1}=\frac{ 1 }{ 2^{1} }=\frac{ 1 }{ 2 }\] \[-4^{2}+2^{-1}=-16+\frac{ 1 }{2^{1} }=-16+\frac{ 1 }{ 2 }=\frac{ -32+1 }{ 2 }=\frac{ -31 }{2 }\]

OpenStudy (anonymous):

\[-16+\frac{ 1 }{2 }=\frac{ -32 }{2 }+\frac{ 1 }{ 2 }=\frac{ -32+1 }{2 }=\frac{ -31 }{ 2 }\]

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

so basically you have to multiply the 16 by 2 and still add 1/2?

OpenStudy (anonymous):

\[\frac{ -16 }{1 }+\frac{ 1 }{2 },L.C.M.of~ denominator~ 1~and~ 2 ~is~2\]

OpenStudy (anonymous):

oh i didn't even think about that